OpenStudy (anonymous):

In a certain federal prison it is known that 2/3 of the inmates are under 25 years of age. It is also known that 3/5 of the inmates are male and that 5/8 of the inmates are female or 25 years of age or older. What is the probability that a prisoner selected at random from this prison is female and at least 25 years old?

OpenStudy (kirbykirby):

Let: \(T\) = inmate under 25 years old \(\overline{T}\) = inmate 25 years or older \(M\) = male \(\overline{M}\)= female You are given: \(P(T)=2/3\) \(P(M)=3/5\) \(P(\overline{M} \cup \overline{T})=5/8\) You need: \(P(\overline{M} \cap \overline{T})\) Thus: \(P(\overline{M} \cup \overline{T})=P(\overline{M})+P(\overline{T})-P(\overline{M} \cap \overline{T})\) \(\implies P(\overline{M}\cap \overline{T})=P(\overline{M})+P(\overline{T})-P(\overline{M}\cup \overline{T})\) \(\implies P(\overline{M}\cap \overline{T})=[1-P(M)]+[1-P(T)]-P(\overline{M}\cup \overline{T})\) \(\implies P(\overline{M}\cap \overline{T})=[1-3/5]+[1-2/3]-5/8\) \(\therefore P(\overline{M} \cap \overline{T})=13/120 \approx 0.1083\)

OpenStudy (anonymous):

THANKS! I was close but somethings weren't matching up in my head.

OpenStudy (kirbykirby):

=]